# A new Angle on Benchmarking Noncovalent Interactions

### v2

@@ Abstract
For noncovalent interactions (NCIs), the CCSD(T) coupled cluster method is widely regarded as the ‘gold standard’. With localized orbital approximations, benchmarks for ever larger NCI complexes are being published; yet tantalizing evidence from quantum Monte Carlo (QMC) results appears to indicate that as the system size grows, CCSD(T) overbinds NCIs by progressively larger amounts, particularly when -stacking is involved. Alas, post-CCSD(T) methods like CCSDT(Q) are cost-prohibitive, which requires us to consider alternative means of estimating post-CCSD(T) contributions.
In this work, we take a step back by considering the evolution of the correlation energy with respect to the number of subunits for such -stacked sequences as acene dimers and alkadiene dimers. We show it to be almost perfectly linear, and propose the slope of the line as a probe for the behavior of a given electron correlation method. By comparison with rank-reduced CCSDT(Q) results for benzene and naphthalene dimers, we show that while CCSD(T) does slightly overbind, it does not at the level suggested by the QMC results.
@@ 1 Introduction
We are trying to glance at a piece of this puzzle by extrapolating dimer interactions to larger and larger sizes using different methods. We will show that by investigating the slope of the intermolecular correlation energies with respect to monomer size, we can estimate the error of benchmark methods for much larger systems even than the ones calculated, gauging their accuracy.
Here, we will focus on CCSD(T) as the gold standard of quantum chemistry and examine all conceivable causes of a possible failure:
@@
The polarizability will increase with the system size of aromatic molecules, and so will the strength of their van der Waals interactions with each other. This polarizability may thus be increasingly hard to calculate with progressively larger basis sets. For example, even small molecules and atoms with huge polarizabilities need many diffuse functions in order to accurately describe these quantities.9
@@
In order to calculate larger systems with CCSD(T), local approximations are utilized. As the number of very small contributions to the interaction energy grows exponentially, their neglect might cumulatively result in substantial errors for larger interacting systems.
@@
While for energies, accurate CCSD(T) values are rather easily computable, CCSD(T) geometries are much scarcer, and there are very few studies actually optimizing intermolecular interaction geometries. This effect is mostly neglected, as fixed geometries are taken into account. For training or benchmarking lower-level methods like density functional theory (DFT), this a rational, or at least expected choice. Neglect of the ’relaxation energy’, however, may conceivably affect comparisons with more accurate methods.
@@
Larger intermolecular interactions are possibly overestimated by the use of perturbative triples in the (T) part, as second-order Møller-Plesset perturbation theory (MP2) strongly overestimates10 intermolecular interactions of aromatic molecules. Various papers have proffered this argument, showing that post-CCSD(T) contributions should become relevant at some point.11, 12, 13 However, one has to bear in mind that while MP2 overestimates interaction energies of stacks, MP3 overcorrects and underestimates (which has inspired the MP2.5 method14). CCSD can be seen as a summation of MP singles, doubles, and disconnected products thereof to infinite order, and underestimates interaction energies as it neglects connected triples (T), which make their first appearance in fourth order. In contrast, however, other post-Hartree-Fock or post-DFT methods such as RPA (random phase approximation15) and DFT-SAPT (density functional theory-based symmetry-adapted perturbation theory16) are argued to yield better, if not correct intermolecular interaction energies of larger dimers.11, 12
@@
For intermolecular interactions with medium to large basis sets, the exact basis set limit is usually bracketed by calculations excluding and including counterpoise corrections,17, 18 with very few notable exceptions, like, for example, the formic acid dimer.19 Thus, an increasingly large gap between these two calculations (including and excluding a basis set correction) would indicate a volatility of a certain post-Hartree-Fock method. Basis set incompleteness errors can also often be estimated using localized orbital or other reduced-scaling approaches.
@@
While the use of CCSD(T) gradients is very often computationally prohibitive, several structures can be considered in order to get a more complete picture of the interaction. This is usually done by just varying the distances, e.g. extending the S66 set20 of intermolecular interactions to the S66x8 set,21 which considers each complex at eight different multiples of the intermonomer separation (at fixed intramonomer geometries), for a total of 528 data points.
@@
Post-Hartree-Fock methods such as coupled cluster theory3 have a clear hierarchy. For example, when static correlation effects are present and not captured well by CCSD(T), one may attempt to walk up the cluster expansion staircase to CCSDT and CCSDT(Q), which will yield some insight about whether ‘the gold standard’ is really that golden.22, 23, 24
@@
The new angle we propose is to compare the behavior of different correlation methods along sequences of successively expanded monomers. We shall show that not only is this behavior rather linear, but that the slopes of the lines, and specifically their deviation from the most accurate method offer a new insight into the (mis)behavior of more approximate correlation methods.
@@ 2 Computational Details
The calculations have been performed with various programs, depending on the method employed. The density functional theory (DFT) 25, 26 geometry optimizations utilizing the -B97M-V 27 functional as well as the DLPNO (Domain-Based Local Pair Natural Orbital)-CCSD(T) 28, 29 calculations have been performed with ORCA 30 , DFT-SAPT31, 32 with MOLPRO33, MP334 with QChem35 because of their RI implementation36, and all dRPA15, 37, LNO (Local Natural Orbitals)-CCSD(T)38, LNO-CCSDT(Q)39, and canonical CCSD(T)4 / CCSDT(Q)40 calculations with the MRCC41 program suite. All LNO-CCSD(T) calculations have been using their respective MP2 basis set corrections as implemented in the MRCC program. While some canonical post-CCSD(T) calculations were also carried out using MRCC, the most demanding ones were performed using a development version of the CFOUR program system42.
For the L7 (i.e., seven large noncovalent complexes) set of systems,43 geometries from the original reference have been utilized as-is. All other structures were optimized with the -B97M-V functional27 using the QZVPP44 basis set with full symmetry. For example, in the case of the benzene dimer, the D6h symmetric saddle point has been utilized. This implies that all structures are directly on top of each other, as only their interaction energy matters. They are not parallely displaced, such as the minimum structures of the benzene45 and coronene dimers46 — they are not minima at all. Despite this, the monomer geometries have either been fixed or relaxed, staying within their respective symmetry for the ethylene stacks, whereas for the acene stacks, we have employed completely relaxed geometries. The rationale for this approach is to be able to use as much symmetry as possible, since only the slope of interaction energies of increasingly large dimers matters.
Throughout this study, Dunning’s correlation consistent basis sets with and without diffuse functions, denoted cc-pVZ47 and aug-cc-pVZ48, have been put to use and will sometimes be abbreviated by DZ, TZ, QZ, and 5Z, as well as aTZ, aQZ, and a5Z. The notations {T,Q}Z and {Q,5}Z refer to basis set extrapolations from TZ and QZ or QZ and 5Z basis sets according to the familiar extrapolation formula, respectively.49
Unless specifically indicated otherwise, interaction energies have been counterpoise (CP) corrected by the standard Boys and Bernardi procedure.50
In order to shed additional light on the performance of CCSD(T)4 vs. fully iterative CCSDT51 and CCSDT(Q),40 we carried out additional calculations using two post-CCSD(T) methods that still only have CPU time scaling with system size N: CCSD(T)Λ 52, 53 and CCSDT-254, 51, 55. Cremer and coworkers56 presented a detailed perturbation theoretical analysis of the differences between full CCSDT and various approximations to it. The CCSDT-3 method omits the coupling term from the full CCSDT amplitudes equations, while CCSDT-2 additionally omits all coupling terms from the said equations. Both reduce the CPU time scaling from to at the expense of an energy difference starting at fifth order in perturbation theory:
@@
where it should be noted56 that is the Hermitian conjugate of .
[…]
@@ 3 Results and Discussion
3.1 L7 Set and the discrepancies between two reference methods
The original starting point of this study was, in fact, the discrepancy between the two reference methods CCSD(T) and FN-DMC (fixed node diffusion Monte Carlo). For this purpose, it is useful to reiterate the differences between them. In addition, we have also performed DFT-SAPT calculations for all these systems. According to the theoretical background of references 11, 12 and point one above, DFT-SAPT may be even more accurate than CCSD(T) for the interaction energies of ever larger systems. In Table 1, we show the DFT-SAPT together with the literature values of various post-Hartree-Fock methods and FN-DMC for four molecules of the L7 set43 as well as the buckycatcher.
@@ 3.1 L7 Set and the discrepancies between two reference methods
| Benzene2 | 10.02 | 7.220 | 6.160 | 11.861 | 11.22 | 13.420 | 19.720, 19.760 |
| GCGC | 52.02 | 36.043 | 35.760 | 54.7 | 57.02 | 56.143 | 74.743, 73.6,62 74.960 |
| C3A | 62.42 | 34.143 | 47.860 | 62.7 | 69.12 | 74.743 | 113.043, 109.2,62 109.760 |
| C2C2PD | 75.62 | 27.743 | 58.260 | 81.5 | 86.12 | 95.443 | 159.943, 154.3,62 155.760 |
| C3GC | 101.32 | 61.843 | 80.860 | 104.6 | 120.02 | 127.243 | 188.843, 182.2,62 183.760 |
| C60CPPA | 130.52 | 93.860 | 157.8 | 174.62 | 354.960 |
@@
It is apparent from these numbers that MP2 often overestimates the interaction energies of large conjugated molecules, whereas MP3 underestimates them. MP2.5,63 which relies on compensation between the errors of MP2 and MP3, still somewhat overestimates the interaction energies when going to larger molecules.64 The values obtained by DFT-SAPT are pretty much exactly between CCSD(T) and fixed-node diffusion Monte-Carlo (FN-DNC) values.2 If we plot interaction energies vs. number of atoms, we obtain values ranging from 0.79 (MP3) to 0.94 (FN-DMC). The slope of FN-DMC is with 1.09 kJ/mol per atom much smaller than the slope of CCSD(T) with 1.46 kJ/mol per atom- with DFT-SAPT and also MP2.5 (1.28 and 1.34 kJ/mol per atom, respectively) in this case being somewhat smaller than CCSD(T).
@@ 3.2 Intermolecular Interaction Energies of Larger Molecules
The first question is: Can we estimate these numbers also from interaction energy slopes of other molecules? In the case that they do, they may offer a hint about the accuracy of different post-Hartree-Fock methods for the interaction energies of large molecules. The systems that may come in mind are the benzene, coronene, and circumcoronene dimers, as the coronene dimer with 72 atoms was already present in the original L7 set. The parallel displaced coronene dimer is yielding a significant difference between FN-DMC with an interaction energy of 75.6 kJ/mol and CCSD(T) with an interaction energy of 86.1 kJ/mol, which is more than 10 kJ/mol. Given, however, that the circumcoronene dimer already has 144 atoms, we need other systems as touchstones for our hypothesis: that we can an interpolate the behavior for really large systems from the slopes of interaction energies of polyaromatic dimers.
To stay as close to the aromatic behavior that we see in the above sequence, we chose the acene dimer series, thus again beginning with the interaction energies of the benzene through hexacene dimers. Starting from heptacene dimer, however, the monomers will undergo cyclodimerization,65, 66, 67 which means that if we want to relax geometries like in the L7 paper (even with symmetry constraints) we will have to stop at hexacene dimer. The latter, with 82 atoms, is about the same size as coronene dimer.
As a second series, we can mix these species, and arrive at the series for the benzene-anthracene, anthracene-pentacene, and so on dimers, whereas in order to retain symmetry, there are the possibility of the CNCN, CNCN-2, CNCN-4 species, and so on – with N denoting the number of rings.
Finally, to perspective, we will also consider at ethylene stacks, as these stacks are the possible smallest conjugated carbon species which can be investigated, such as the dimers of ethylene, trans-butadiene, hexatriene, octatetraene, and so on. All these species are interesting in the context of quantum chemistry due to their extended -conjugation. As the chain length increases, the extent of -conjugation also increases, resulting in enhanced stability and stronger interactions between the molecules. For larger systems, pure London forces (as distinct from -stacking) become significant for these dimers as well.
@@ 3.3 Basis Set Effects
First, we try to evaluate the basis set effects of the different methods. If, for example, polarization would be a major issue, the dissociation energies along a chain dimer sequence would not be on a straight line, but rather be flat, as the increasing polarization would be less well described by a small number of basis functions. The number of basis functions, however, also increases linearly with system size, yielding almost perfectly straight lines for different acene series, as shown in Figure 1a-c for LNO-CCSD(T) using tight settings. As counterpoise corrections for LNO and similar methods in small molecules are problematic owing to screening out a substantial part of the ”ghost atoms” contribution [see, e.g., Ref.68], we had to exclude the first couple of cp-corrected interaction energies (for example, the benzene and anthracene dimer for the DZ basis set and the benzene, anthracene and naphthalene dimer for the TZ basis set) in each graph.
@@
As it becomes obvious, all of the investigated series are on almost perfectly straight lines with an value of 0.992 for all basis sets. With increasing basis set size, the lines converge nicely to slopes of 24.930.23 kJ/subunit for the acene stacks. Here, the cp-corrected slopes of the aug-cc-pVQZ, cc-pV5Z, and aug-cc-pV5Z basis sets, as well as the non-cp-corrected slopes cc-pVQZ, cc-pVQZ, and aug-cc-pV5Z basis sets fall within this range. Interestingly, while the intercepts of these graphs seem highly basis set dependent, the slopes are much less sensitive to the basis set used. For example, a basis set of aug-cc-pVTZ quality deviates in the intercept by 33% from the aug-cc-pV5Z basis set for the cp-corrected correlation energy of the acene series, but in its slope by only 3%.
For the ethylene stacks, the intercept of cc-pVTZ deviates from aug-cc-pV5Z sets by 13%, and the slope by only 3%. This trend is visible for all the systems, which implies that even quite small basis sets may give an indication about the slope (albeit not the intercept) of the correlation energy and its behavior. When comparing ethylene to acene stacks, we get about a third of the slopes of the rings (acenes) per double bond (ethylene stacks). This shows that acene stacks and ethylene stacks are still somewhat related. In contrast, the intercept of the ethylene stacks is about three times the intercept of the acene stacks. The ethylene stack series has a slope of 8.700.27 kJ if we consider all basis sets except cc-pVDZ. Based on the obtained graphs and considering basis set sizes, we conclude that utilizing the cc-pVQZ basis set is probably most advantageous for obtaining reliable slopes for both series.

### v3

@@ Abstract
For noncovalent interactions, the CCSD(T) coupled cluster method is widely regarded as the ‘gold standard’. With localized orbital approximations, benchmarks for ever larger complexes are being published; yet tantalizing evidence from FN-DMC (fixed node quantum Monte Carlo) results appears to indicate that as the system size grows, CCSD(T) overbinds intermolecular interactions to a progressively larger degree, particularly when -stacking is involved. Alas, post-CCSD(T) methods like CCSDT(Q) are cost-prohibitive, which requires us to consider alternative means of estimating post-CCSD(T) contributions.
In this work, we take a step back by considering the evolution of the correlation energy with respect to the number of subunits for such -stacked sequences as acene dimers and alkadiene dimers. We show it to be almost perfectly linear, and propose the slope of the line as a probe for the behavior of a given electron correlation method. By comparison with rank-reduced CCSDT(Q) results for benzene and naphthalene dimers, we show that while CCSD(T) does slightly overbind, it does not reach the level suggested by the FN-DMC results.
@@ 1 Introduction
We are trying to glance at a piece of this puzzle by extrapolating dimer interactions to progressively larger scales using different methods. We will show that by investigating the slope of the intermolecular correlation energies with respect to monomer size, we can estimate the error of benchmark methods for much larger systems even than those calculated, gauging their accuracy.
Here, we will focus on CCSD(T) as the gold standard of quantum chemistry and examine all conceivable causes of its possible failure:
@@
Monomer polarizabilities will increase with the system size of aromatic molecules, as will the strength of their van der Waals interactions. This polarizability may thus become increasingly hard to calculate, requiring progressively larger basis sets. For instance, even small molecules and atoms with large polarizabilities require many diffuse functions in order to accurately describe these quantities.9
@@
For calculating larger systems with CCSD(T), local orbital approximations are utilized. As the number of very small contributions to the interaction energy grows exponentially, their collective neglect might conceivably result in substantial errors for larger interacting systems. (“Many a little makes a mickle”, as the English proverb goes.)
@@
While accurate CCSD(T) energies are comparatively easy to compute, CCSD(T) geometries are much scarcer, and few studies actually consider optimized intermolecular interaction geometries. This effect is mostly neglected, as fixed geometries are taken into account. For training or benchmarking lower-level methods like density functional theory (DFT), this a rational, or at least expected choice. Neglect of the ’relaxation energy’, however, may conceivably affect comparisons with more accurate methods.
@@
Larger intermolecular interactions may possibly overestimated due to the quasiperturbative triples in the (T) part, as second-order Møller-Plesset perturbation theory (MP2) is known (e.g., Ref.10) to strongly overestimate10 intermolecular interactions of aromatic molecules. Various papers have put forward this argument, showing that post-CCSD(T) contributions should become relevant at some point.11, 12, 13 However, one has to bear in mind that while MP2 overestimates interaction energies of stacks, MP3 overcorrects, and hence underbinds them (which has inspired the MP2.5 method14). CCSD can be seen as a summation of MP singles, doubles, and their disconnected products to infinite order, and underestimates interaction energies as it neglects connected triples (T), which first appear at fourth order. In contrast, however, other post-Hartree-Fock or post-DFT methods such as RPA (random phase approximation15) and DFT-SAPT (density functional theory-based symmetry-adapted perturbation theory16) are argued to yield better intermolecular interaction energies for larger dimers.11, 12
@@
For intermolecular interactions with medium to large basis sets, the exact basis set limit is usually bracketed by calculations excluding and including counterpoise corrections,17, 18 with very few notable exceptions, like, for example, the formic acid dimer.19 Thus, an increasingly large gap between these two calculations (including and excluding a basis set correction) would indicate a volatility of a certain post-Hartree-Fock method. Basis set incompleteness errors can often also be estimated using localized orbital or other reduced-scaling approaches.
@@
While the use of CCSD(T) gradients is very often computationally prohibitive, several structures can be considered to obtain a more complete picture of the interaction. This is usually done by just varying the distances, e.g. extending the S66 set20 of intermolecular interactions to the S66x8 set,21 which considers each complex at eight different multiples of the intermonomer separation (with fixed intramonomer geometries), for a total of 528 data points.
@@
Post-Hartree-Fock methods such as coupled cluster theory3 have a clear hierarchy. For example, when static correlation effects are present and not well-captured by CCSD(T), one may attempt to walk up the cluster expansion staircase to CCSDT and CCSDT(Q), which will yield some insight about whether ‘the gold standard’ is really that golden.22, 23, 24
@@
The new angle we propose is to compare the behavior of different correlation methods along sequences of successively expanded monomers. We shall show that, not only is this behavior rather linear, but that the slopes of the lines — specifically their deviation from the most accurate method — offer a new insight into the (mis)behavior of more approximate correlation methods.
@@ 2 Computational Details
The calculations have been performed with various programs, depending on the method employed. The DFT25, 26 geometry optimizations utilizing the -B97M-V 27 functional as well as the DLPNO (Domain-Based Local Pair Natural Orbital)-CCSD(T) 28, 29 calculations have been performed with ORCA 30 , DFT-SAPT31, 32 with MOLPRO33, MP334 with QChem35 because of their RI implementation36, and all dRPA15, 37, LNO (Local Natural Orbitals)-CCSD(T)38, 39, LNO-CCSDT(Q)40, and canonical CCSD(T)4, 41 / CCSDT(Q)42 calculations with the MRCC43 program suite. All LNO-CCSD(T) calculations have been using their respective MP2 basis set corrections as implemented in the MRCC program. While some canonical post-CCSD(T) calculations were also carried out using MRCC, the most demanding ones were performed using a development version of the CFOUR program system44. The DFT-SAPT calculations have been performed using the asymptotically corrected PBE0 functional 45, 46, 47, 48, where the monomer ionization potentials and energies of the highest occupied molecular orbital have been computed via PBE0 calculations using a cc-pVQZ basis set.
For the L7 (i.e., seven large noncovalent complexes) set of systems49 and the buckycatcher,2 geometries from the original references have been utilized as-is. All other structures were optimized with the -B97M-V functional27 using the QZVPP50 basis set with full symmetry. For example, in the case of the benzene dimer, the symmetric saddle point has been utilized. This implies that all structures are directly on top of each other, as only their interaction energy matters. They are not parallel-displaced, such as the minimum-energy structures of the benzene51 and coronene dimers52 — they are not minima at all. The monomer distances have either been fixed, staying within their respective distance for the ethylene stacks, or completely relaxed within their symmetry for both the ethylene as well as for the acene stacks. The rationale for this approach is to be able to use as much symmetry as possible, since only the slope of interaction energies of increasingly large dimers matters.
Throughout this study, Dunning’s correlation consistent basis sets with and without diffuse functions, denoted cc-pVZ53 and aug-cc-pVZ54, have been put to use and will sometimes be abbreviated by DZ, TZ, QZ, and 5Z, as well as aTZ, aQZ, and a5Z. The notations {T,Q}Z and {Q,5}Z refer to basis set extrapolations from TZ and QZ or QZ and 5Z basis sets according to the familiar extrapolation formula, respectively.55
Unless specifically indicated otherwise, interaction energies have been counterpoise (CP) corrected by the standard Boys and Bernardi procedure.56
@@
where is the perturbation parameter and is the fifth-order triples-triples interaction term. The additional terms CCSDT-2 carries beyond CCSD(T) again start out at fifth order:
@@
where it should be noted60 that is the Hermitian conjugate of .
@@ 3 Results and Discussion
3.1 L7 Set and discrepancies between fixed-node diffusion Monte Carlo and CCSD(T)
The original starting point of this study was, in fact, the discrepancy between the two reference methods CCSD(T) and FN-DMC (fixed node diffusion Monte Carlo). For this purpose, it is useful to reiterate the differences between them. In addition, we have also performed DFT-SAPT calculations for all these systems. The theoretical background sections of Refs.11, 12 argue that DFT-SAPT may be even more accurate than CCSD(T) for the interaction energies of progressively larger systems. In Table 1, we show the DFT-SAPT together with the literature values of various post-Hartree-Fock methods and FN-DMC for four molecules of the L7 set49 as well as the buckycatcher.
@@ 3.1 L7 Set and discrepancies between fixed-node diffusion Monte Carlo and CCSD(T)
| Benzene2 | 10.02 | 7.220 | 6.162 | 11.863 | 11.22 | 13.420 | 19.720, 19.762 |
| GCGC | 52.02 | 36.049 | 35.762 | 54.7 | 57.02 | 56.149 | 74.749, 73.6,64 74.962 |
| C3A | 62.42 | 34.149 | 47.862 | 62.7 | 69.12 | 74.749 | 113.049, 109.2,64 109.762 |
| C2C2PD | 75.62 | 27.749 | 58.262 | 81.5 | 86.12 | 95.449 | 159.949, 154.3,64 155.762 |
| C3GC | 101.32 | 61.849 | 80.862 | 104.6 | 120.02 | 127.249 | 188.849, 182.2,64 183.762 |
| C60CPPA | 130.52 | 93.862 | 157.8 | 174.62 | 354.962 |
@@
It is apparent from these numbers that MP2 often overestimates the interaction energies of large conjugated molecules, whereas MP3 underestimates them. MP2.5,65 which relies on compensation between the errors of MP2 and MP3, still somewhat overestimates the interaction energies when going to larger molecules.66 The values obtained by DFT-SAPT are pretty much exactly between CCSD(T) and fixed-node diffusion Monte-Carlo (FN-DNC) values.2 If we plot interaction energies vs. number of atoms, we obtain coefficients of determination ranging from 0.79 (MP3) to 0.94 (FN-DMC). The slope of FN-DMC, at 1.09 kJ/mol per atom, is much smaller than that of CCSD(T) at 1.46 kJ/mol per atom — with DFT-SAPT and MP2.5 (1.28 and 1.34 kJ/mol per atom, respectively) both yielding somewhat smaller slopes than CCSD(T).
@@ 3.2 Intermolecular Interaction Energies of Larger Molecules
The first question is: Can we also estimate these numbers from interaction energy slopes of other molecules? If so, they may offer a hint about the accuracy of different post-Hartree-Fock methods for the interaction energies of large molecules. Systems that come to mind are the benzene, coronene, and circumcoronene dimers, as the coronene dimer with 72 atoms was already present in the original L7 set. The parallel-displaced coronene dimer exhibits a significant difference in excess of 10 kJ/mol between FN-DMC, 75.6 kJ/mol, and CCSD(T), 86.1 kJ/mol. Given, however, that the circumcoronene dimer already has 144 atoms, we need other, more tractable, systems as touchstones for our hypothesis, hopefully enabling us to infer the behavior for really large systems from the slopes of interaction energies of polyaromatic dimers.
To closely match the aromatic behavior observed in the above sequence, we chose the acene dimer series, thus again starting with the interaction energies of the benzene through hexacene dimers. Starting from heptacene dimer, however, the monomers will undergo cyclodimerization,67, 68, 69 which means that if we aim to relax geometries like in the L7 paper (even with symmetry constraints) we will have to stop at hexacene dimer. The latter, with 82 atoms, is about the same size as coronene dimer.
As a second series, we can mix-and-match species, and arrive at the series including benzene-anthracene, naphthalene-tetracene, anthracene-pentacene, and pentacene-heptacene. In order to retain symmetry of the pi-stacks directly on top of each other, we can construct lines using the CNCN, CNCN-2, and CNCN-4 series (with N denoting the number of rings ranging from one to six/seven).
Finally, for perspective, we will also consider ethylene stacks, as these stacks are the possible smallest possible conjugated carbon species which can be investigated, such as the dimers of ethylene, trans-butadiene, hexatriene, octatetraene, and so on. All these species are interesting in the context of quantum chemistry due to their extended -conjugation. As the chain length increases, the extent of -conjugation also increases, resulting in enhanced stability and stronger interactions between the molecules. For larger systems, pure London dispersion forces (as distinct from -stacking) become significant for these dimers as well.
@@ 3.3 Basis Set Effects
First, we try to evaluate the basis set effects of the different methods. If, for example, polarization were a major issue, the dissociation energies along a chain dimer sequence would not be on a straight line, but rather taper off, as the increasing polarization would be increasingly poorly described by a limited number of basis functions. The number of basis functions, however, also increases linearly with system size, yielding almost perfectly straight lines for different acene series, as shown in Figure 1a-c for LNO-CCSD(T) using tight settings. As counterpoise corrections for LNO and similar methods in small molecules are problematic because a substantial part of the ”ghost atoms” contribution is getting screened out [see, e.g., Ref.70], we had to exclude the first couple of counterpoise-corrected interaction energies (for example, benzene and naphthalene dimers for the DZ basis set, and the benzene, naphthalene, and anthracene dimers for the TZ basis set) in each graph.
@@
As becomes apparent, all of the investigated series are on almost perfectly straight lines with an for all basis sets. With increasing basis set size, the lines converge nicely to slopes of 24.930.23 kJ per subunit for the acene stacks. Here, the CP-corrected slopes of the aQZ, 5Z, and a5Z basis sets, as well as the non-CP-corrected slopes QZ, aQZ, and a5Z basis sets fall within this range. Interestingly, while the intercepts of these graphs seem highly basis set-dependent, the slopes are much less sensitive to the basis set used. For example, for the CP-corrected correlation energy of the acene series, the intercepts for basis sets of aTZ and a5Z quality deviate by about one-third, but the slopes only by about 3%.
For the ethylene stacks, the intercept for TZ deviates from that for a5Z by 13%, but (again) the slope by only 3%. This trend is visible across all systems, suggesting that even quite small basis sets may give an indication about the slope (albeit not the intercept) of the correlation energy and its behavior. When comparing slopes for ethylene vs. acene stacks, we find the former to be roughly one-third of the latter. This suggests that there is still some relationship between their behaviors. In contrast, the intercept of the ethylene stacks is about three times the intercept of the acene stacks. The ethylene stack series has a slope of 8.700.27 kJ if we consider all basis sets except DZ. Based on the obtained graphs and considering basis set sizes, we conclude that the QZ basis set is likely the most advantageous for obtaining reliable slopes for both series.